what is the value of x if x + 60 is supplementary to x + 10?
Answer by ankor@dixie-net.com(15622)

what is the value of x if x + 60 is supplementary to x + 10?
:
Two angles that are supplementary, add up to 180, therefore
(x+60) + (x+10) = 180
2x + 70 = 180
2x = 180 - 70
2x = 110
x = 110/2
x = 55